two types of milk costing rs 20 per litre and rs 24 per litre are mixed in equal volume. then water is added making the mixture of 66 litres. if it is sold at the rate of rs 16 per litre,what is the ratio of milk and water in the mixture ?
Answer by Edwin McCravy(20055)

two types of milk costing rs 20 per litre and rs 24 per litre are mixed in equal volume. then water is added making the mixture of 66 litres. if it is sold at the rate of rs 16 per litre,what is the ratio of milk and water in the mixture ?
We mix 
x litres of low cost milk
x litres of high cost milk
y litres of water

x + x + y = 66    <--- liquid equation
   2x + y = 66    

Price of x litres of low cost milk = 20x rs
Price of x litres of high cost milk = 24x rs
Price of y litres of water = 0y = 0 rs
Price of all 66 litres = 66×16 = 1056 rs

20x + 24x + 0 = 1056   <--- rupees equation
          44x = 1056   
            x = 24 liters of each type milk.

Substitute in 2x + y = 66
           2(24) + y = 66
              48 + y = 66
                   y = 18 litres of water

Amount of milk = 24 litres (low cost) + 24 litres (high cost) = 48 litres milk
Amount of water = 18 litres

Ratio of milk to water = 48:18 = 8:3

Edwin